The combined equation to a pair of straight lines pass through the origin and inclined 30degrees and 60degrees respectively with x-axis
 

Grade:8

2 Answers

Arun
4 years ago
Dear student
 
First find out seperate equation
 
y = √3 x
y - √3 x = 0
And 
y =x /√3
 
y - x/√3 = 0
 
Now for combined equation
 
Multiply
 
(y - √3 x) ( y - x/√3) = 0
 
y² +x² -xy (√3 + (1/√3))  = 0
Vikas TU
3 years ago
We need to find two straight line and the product of the two will give pair of straight lines equation
suppose 1st =y=mx+c  passing through origin and slope is tan30
2nd y=mx+c passing through origin and slope is tan60